A piece of string is 122 inches long if u cut it into 2 pieces so that one piece is three times as long as the other how long is each piece?
Answer by math_helper(2461)


x + 3x = 122
4x = 122
x = 122/4 = 30.5 in
---
One piece is 30.5 in, the other is 91.5 in